Please help me solve this problem: How much of a 50% acid solution must be mixed with 30 gallons of an 18% acid solution to obtain a solution that is 40% acid?
Answer by teentutor(5)

My guess is that your teacher taught you the 3 by 3 grid system for figuring out these kinds of problems, so my tutoring will be based on that. If not, you can look it up on the internet and know exactly what I am talking about. So, plug in the percentages into the top row, so it will look like 50% l 18% l 40%. Then, in the second row, plug in the amounts, using x and y, so x l 30 l y . Next, you can multiply to a decimal and variable. .5x l 5.4 l .4y. Now, use substitution. it will look like x+30=y
.5x-5.4=.4y Multiply so that both equation come out with opposite variables to cancel out, in out case -2. now you will get
x+30 =y
-x+10.8=-.8y
Simplify to 40.8 = .2y. So y=204. since x+30=204, x=174.